# Deep-learning-based continuous attacks on QKD protocols

### v2

@@ Abstract
The most important characteristic of a Quantum Key Distribution (QKD) protocol is its security against third-party attacks, and the potential countermeasures available. While new types of attacks are regularly developed in the literature, they rarely involve the use of weak continuous measurement. Here, we design a new attack scheme called Deep-learning-based continuous attack (DLCA) that exploits continuous measurement together with the powerful pattern recognition capacities of deep recurrent neural networks. We show that, when applied to the BB84 protocol, our attack can be difficult to notice while still allowing the spy to extract significant information about the states of the qubits sent in the quantum communication channel. Finally, we study how the spy can exploit quantum feedback to further cover their tracks. Our attack scheme, while still at the early stages of a toy model, constitutes a potential threat which is worthwhile to be investigated, also as it could be applied to different QKD protocols and generalized in many different ways.
@@ I Introduction
At the same time, quantum computing represents also a threat in terms of security, in particular related to some asymmetric cryptographic algorithms such as RSA (Rivest-Shamir-Adleman) [3], a public key cryptosystem still used in many secure data transmissions to this day. Indeed, standard encryption techniques such as RSA, while very powerful, could be broken through Shor’s algorithm, a quantum algorithm factoring large integers exponentially faster than the best-known classical algorithms [4].
@@
While current quantum computing technology is still far from being enough advanced to break RSA, this motivated the elaboration of new encryption techniques based on quantum mechanical properties. Quantum Key Distribution (QKD), which aim is to implement the exchange of a private key over a public insecure channel between two parties, is the most famous category of quantum cryptography protocols [5]. The first and most known QKD protocol is the BB84, proposed by Charles Bennett and Gilles Brassard in 1984 [6], and uses linearly polarized photons travelling in an optical fiber. Among others are the B92 that uses entangled particles [7], the Differential-phase-shift which does not require a basis selection [8] or the Decoy State protocol designed to overcome photon number splitting attacks [9].
@@
One of the main motivations to look for quantum cryptography protocols over classical ones originates from the perturbative nature of measurement in quantum mechanics. Indeed, any spy acting on a communication channel will influence the states of the qubits used to store the private key bits traveling inside, because of the collapse of the wavefunction, which makes the spy more easily detectable than in classical communication protocols. Amongst the most studied types of attacks are the Intercept-and-Resend type, Photon Number Splitting (PNS) [10] and Trojan Horse (also called Large Pulse attack) [11, 12]. New Intercept-and-Resend attacks have been developed recently, such as Blinding [13], Time shift [14, 15] or Dead-time [16]. While some attacks employ strategies to make the spy measurement effect less detectable [e.g., by thermally blinding Bob’s detector [13] or by monitoring the electromagnetic emissions of a QKD emitter and using a deep convolutional neural network (Deep-learning-based radio-frequency side-channel attack [17])], several of these attacks include projective measurements that usually provoke a perturbation on the photon state. Hence, if the two parties, say Alice and Bob, do not deploy specific countermeasures against the attacks [18], they can usually find a way to decide or not on the presence of a spy on the quantum communication channel, as studied in the past, although not extensively [19, 20, 21, 22]. In this respect, a typical measure Alice and Bob can calculate (by sacrificing a few bits of the private key by sharing their measurement results on a public channel) is the Quantum Bit Error Rate (QBER), which is defined as the rate of incorrect results Bob gets when measuring the qubits sent by Alice in the right basis [23], i.e.,
@@
where $N_{\mathrm{total}}$ is the total number of qubits received where Bob used the right measurement basis, and $N_{\mathrm{error}}$ is the number of incorrect results he gets among these qubits. In the case of a perfect quantum communication channel, the QBER should be zero. However, in the presence of a spy, the qubits states are usually altered and the QBER non-zero despite Alice and Bob use the same basis for the qubits, which should thus signal the two parties that something went wrong. Complications then arise because in practice, the quantum channel is not perfectly isolated from its environment (the optical fiber could be leaky [24, 25, 26]) and the measurement apparatus of Bob could be defective, which contribute to another cause of QBER enhancement. Hence, distinguishing an attack from an intrinsic noise in the channel is not always easy.
@@
In this paper, we develop a new type of attack based on continuous (homodyne) measurement [27] of individual polarized photons and apply it for concreteness in the context of the BB84 protocol. The general motivation beyond this is to evaluate the performance of attacks that produce only a small perturbation to the qubits, by contrast with the effects of projective measurement usually involved in the other types of attacks. In particular, we investigate how the effects of the spy measurement can be optimally hidden by the intrinsic noise of the quantum communication channel by minimizing the increase of the QBER due to the measurement. In the same time, we study how to maximize the information gain by the spy by feeding the outcome of the continuous measurement, also called homodyne photocurrents, to a Long Short-Term Memory (LSTM) recurrent neural network [28] to retrieve the initial states of the photons sent by Alice, which compose the private key generated.
@@
This paper is organized as follows: In Sec. II we first remind the BB84 protocol step by step, present our model of the qubit dynamics in the quantum communication channel when subject to intrinsic dissipation and continuous measurement, investigate how a spy could use the outcome of this measurement to obtain the initial state of the qubit and present the neural network we implemented to do so. In Sec. III we compare the results obtained via a basic projective measurement (Intercept-and-Resend attack) and our measurement scheme. We then extend the theory to include quantum feedback, to try to decrease the measurement impact while keeping the same extracted information. Finally, in Sec. IV, we conclude and discuss potential perspectives.
@@ II Model and methods
In this section, we first remind how the standard BB84 protocol works briefly, before presenting how we model the dynamics of the qubits used in the protocol when they are subjected to dissipation and continuous measurement in the quantum communication channel. Then, we describe the neural network that we envision a spy could use to retrieve the states of the qubits based on the continuous measurement they performed in the channel. Finally, we present how we quantify the impact of the measurement on the security of the protocol.
@@ 5. Step 5:
    Alice selects a subset of her bits to check on the interference caused by the spy – the so-called Eve –, and tells Bob which bits she chose. They both announce and compare the values of the check bits via the public channel and calculate the QBER given by Eq. (1). If it is higher than a threshold, they abort the protocol.
To increase the security of the protocol, two additional steps could be performed: information reconciliation (i.e., correcting bits that have been modified by dissipation [30, 31]), or privacy amplification (i.e., passing the generated key in a hash function to decrease Eve mutual information [30, 32, 33]), but which are not the purpose of this paper.
@@ II.2 Dissipative qubit dynamics conditioned on measurement
Figure 1: Sketch representation of the attack scheme we develop in this paper applied on the BB84 protocol. Alice (A) sends linearly polarized photons to Bob (B) via an optical fiber, while travelling they are subject to dissipation with a rate $\gamma_{D}$ and to weak measurement with a rate $\gamma_{E}$, as could be done e.g. via monitoring of a field coupled to the photon states of the optical fiber. The output of this measurement is treated using a neural network to determine the initial state. In the sketched example, Alice sends six photons to Bob, among which three are polarized in the Pauli-X eigenbasis and three in the Pauli-Z eigenbasis. Bob measures randomly in one of these two bases each of the photons received, and the basis chosen is the right one for four of these. However, even if the measurement basis is right, the two results circled in red are wrong due to dissipation and measurement that changed the qubit state in the optical fiber. In this example, if Alice and Bob compare all the measurements when they chose the same basis, the QBER [Eq. (1)] would be $50\%$.
@@
$+\gamma_{E}\mathcal{D}[e]\rho_{J}dt+\sqrt{\gamma_{E}\eta}dW\mathcal{H}[e]\rho_{J},$
where $H$ is the channel Hamiltonian defined as $H=\omega\sigma_{z}$ for the initial states |$\ket{0}$ or |$\ket{1}$ and $H=\omega\sigma_{x}$ for $\ket{+}$ or $\ket{-}$ [24] with $\sigma_{x}=|0\rangle\langle 1|+|1\rangle\langle 0|$ and $\sigma_{z}=|0\rangle\langle 0|-|1\rangle\langle 1|$ the standard Pauli operators, where $\rho_{J}$ is the density operator of the qubit conditioned on the measurement with efficiency $\eta\in[0,1]$ of the homodyne current
@@
to model the dissipation as a bit-flip error, but any other choice could be made without any additional complexity, depending e.g. on the specific optical fiber open system model investigated.
@@ II.4 Neural network quantum tomography based on the measurement
The input layer of our network is a LSTM one with $40$ units in its hidden state, to take as input the time series data that are the photocurrents. We then use two dense hidden layers of $40$ and $20$ neurons, with activation functions set to ReLU and sigmoid respectively. The output layer is a $4$ neurons softmax layer. The loss function we use for training is the sparse categorical cross entropy since we deal with a $4$-class classification problem, and the optimizer is Adam with default parameters. The training consists in one epoch over $9\times 10^{4}$ photocurrents while the testing is done on $10^{4}$ photocurrents. Before entering the network the currents are standardized and flipped along their time axis.
@@
II.5 Impact of measurement VS Information gain
As explained earlier, the goal of the spy is to minimize their impact on the photon states while maximizing their information gain about the initial state sent by Alice. To quantify the impact of the measurement, we use the QBER introduced earlier [Eq. (1)], as it is directly measurable by Alice and Bob and allow them to assess the security of the protocol. To quantify the success of the spy in retrieving the initial state of the photons, we use the spy accuracy that we will denote by $A$. In the context of our deep learning approach, this is the neural network test accuracy [45] which is defined as the percentage of good predictions among all the predictions of the network on the test set. However, when dealing with a projective measurement, this accuracy is defined as the probability that the spy measures the right state, as elaborated on below.
@@ III Results
In this section, we study the impact of our attack and the information gained by the spy in different cases, from the analysis of the QBER and the accuracy $A$. We first compute the QBER in the case of no attack. Then, we study a simple standard projective measurement attack, before investigating our continuous measurement scheme. Finally, we close the section by studying the performance of a quantum feedback based on the measurement the spy could use to cover their tracks.
@@ III.1 No attack
In the case where no spying is done on the quantum channel, which means there is no measurement and only the intrinsic dissipation, the QBER can easily be obtained from Eq. (3) with $e=0$, which corresponds to a Lindblad master equation (see Appendix A), and reads
@@ III.2 Attack via projective measurement
In this case, the information gain by the spy, or equivalently the accuracy $A$, can be calculated exactly by solving the Lindblad master equation with a single jump operator $L=\sigma_{x}$ (see Appendix B) and reads
@@
which depends on the time $t^{*}$ at which the projective measurement is performed. Therefore, Eve must measure the photons as close to Alice as possible in order to maximize Eq. (9) and get as much information as possible on the private key, which is here bounded by $75\%$, meaning that the spy has at best $75\%$ chance to guess the initial state sent by Alice.
The QBER can also be obtained easily (see Appendix C), and reads
@@
Interestingly, we see that the time $t^{*}$ at which Eve performs her measurement does not impact the probability that Bob measures the state he is supposed to. Also, comparing Eqs. (8) and (10), we clearly see that Alice and Bob will always be able to distinguish the presence of the spy from intrinsic dissipation. This result is displayed as the dotted black line in Fig 3.
Figure 3: Top: QBER as a function of $\gamma_{D}t$ for different attack schemes and angles $\theta$. The black dashed line represents the QBER when no measurement is performed (i.e., produced only by intrinsic dissipation) [Eq. (8)]. The dotted black line is the QBER with dissipation and projective measurement [Eq. (10)] and the red line is the QBER caused by a continuous measurement with $e=\sigma_{z}$ ($\theta=\pi/2$). The orange line is the QBER when using the angle $\theta_{A}^{*}$ maximizing the accuracy (which we update every $\gamma_{D}t=0.3$), the green line for the angle $\theta_{qber}^{*}$ minimizing the QBER, and the blue line for the angle $\theta_{\lambda}^{*}$ minimizing $\lambda$. Bottom: optimized angles with respect to the QBER itself (green line), the NN test accuracy (orange line) and $\lambda$ (blue line).

### v3

@@ Abstract
The most important characteristic of a Quantum Key Distribution (QKD) protocol is its security against third-party attacks, and the potential countermeasures available. While new types of attacks are regularly developed in the literature, they rarely involve the use of weak continuous measurement and more specifically machine learning to infer the qubit states. In this paper, we design a new individual attack scheme called Deep-learning-based continuous attack (DLCA) that exploits continuous measurement together with the powerful pattern recognition capacities of deep recurrent neural networks. We show that, when applied to the BB84 protocol, our attack increases only slightly the Quantum Bit Error Rate (QBER) of a noisy channel and allows the spy to infer a significant part of the sifted key. In addition, we show it yields similar performances in terms of information gain when compared to an optimal individual attack, namely the phase-covariant quantum cloner. Our individual attack scheme demonstrates deep-learning-enhanced quantum state tomography applied to QKD and could be generalized in many different ways,
@@ I Introduction
At the same time, quantum computing also represents a threat in terms of security, in particular related to some asymmetric cryptographic algorithms such as RSA (Rivest-Shamir-Adleman) [3], a public key cryptosystem still used in many secure data transmissions to this day. Indeed, standard encryption techniques such as RSA could be broken through Shor’s algorithm, a quantum algorithm factoring large integers exponentially faster than the best-known classical algorithms [4].
@@
While current quantum computing technology is still far from being enough advanced to break RSA, this motivated the elaboration of encryption techniques based on quantum mechanical properties. Quantum Key Distribution (QKD), which aim is to implement the exchange of a secure private key over a public insecure channel between two parties, is the most famous category of quantum cryptography protocols [5]. The first and most known QKD protocol is the BB84, proposed by Charles Bennett and Gilles Brassard in 1984 [6], which uses linearly polarized photons traveling in an optical fiber. Among others are the B92 that uses entangled particles [7], the Differential-phase-shift which does not require a basis selection [8] or the Decoy State protocol designed to overcome photon number splitting attacks [9].
One of the main motivations to look for quantum cryptography protocols over classical ones originates from the perturbative nature of measurement in quantum mechanics. Indeed, any spy acting on a communication channel will influence the states of the qubits used to store the private key bits traveling inside, because of the collapse of the wavefunction, which makes the spy more easily detectable than in classical communication protocols. In other words, any attempt to retrieve information from the system will inevitably introduce some disturbance, described by the Information-Disturbance theorem [10, 11].
@@
Amongst the most studied types of attacks are the Intercept-and-Resend type, Photon Number Splitting (PNS) [12] and Trojan Horse (also called Large Pulse attack) [13, 14]. New Intercept-and-Resend attacks have been developed recently, such as Blinding [15], Time shift [16, 17] or Dead-time [18]. While these attacks fall under the category of quantum hacking—exploiting vulnerabilities in practical implementations—many involve projective measurements that typically disturb the quantum state of the photon. For a comprehensive review, we refer the reader to Xu. et al. [19]. If the two parties, say Alice and Bob, do not deploy specific countermeasures against the attacks [20], they can however usually find a way to decide or not on the presence of a spy on the quantum communication channel [21, 22, 23, 24], by sacrificing a few bits of the sifted key (i.e., by sharing their measurement results on a public channel) and calculating the Quantum Bit Error Rate (QBER), which is defined as the rate of incorrect results Bob gets when measuring the qubits sent by Alice in the right basis [25], i.e.,
@@
where $N_{\mathrm{total}}$ is the total number of qubits received where Bob used the right measurement basis, and $N_{\mathrm{error}}$ is the number of incorrect results he gets among these qubits. In the case of a perfect quantum communication channel, the QBER should be zero. However, in the presence of a spy, the qubits states are usually altered and the QBER non-zero despite Alice and Bob using the same basis for the qubits, which should thus signal the two parties that something went wrong. Complications then arise because in practice, the quantum channel is not perfectly isolated from its environment (e.g., the optical fiber could be leaky [26, 27, 28]) and Bob measurement apparatus could be defective, contributing to another cause of QBER enhancement. Hence, distinguishing an attack from intrinsic noise in the channel is not always easy. Despite this, several fundamental papers established the unconditional security of the BB84 protocol in the early 2000s [29, 30, 5, 31, 32]. The security is unconditional in the sense that no assumption is made on Eve’s attacks: Eve can perform any measurement scheme on the channel and the channel may be subject to dissipation, Alice and Bob will upper-bound Eve’s obtainable information and, provided the QBER is below a certain threshold, reduce it to an arbitrary low level using privacy amplication.
The recent development of Machine Learning (ML) and Deep Learning (DL) techniques has led to many improvements in QKD, whether to enhance existing protocols or to detect attacks more easily. Indeed, DL has been used to identify if an attacker is present or not in an IoT network, based on the final key length [33]. In Continuous-Variable QKD (CV-QKD), ML has been used for wavelength-attack recognition [34] and calibration-attack recognition [35]. A single neural network was trained to detect calibration-attacks, LO-intensity-attacks and saturation-attacks, or two types of hybrid attack strategies [36]. Tunc et al. implemented a recurrent neural network and a support vector machines algorithm to protect the BB84 protocol against attacks [37]. Such tools have also been implemented in CV-QKD protocols for, among other things, noise filtering [38], wavefront correction [39], state classification [40], parameter estimation [41] and parameter optimization [42].
However, only a couple of works have investigated how artificial intelligence could be used to develop more effective attacks: a convolutional neural network was trained to help the eavesdropper choose the best opportunity to launch an entanglement-distillation-attack [43], an ML algorithm was shown to be able to analyze the power originating from the integrated electrical control circuit to perform a power-analysis-attack [44], and a quantum circuit implementation of the BB84 protocol was interpreted as a quantum machine learning task, allowing to find a cloning algorithm outperforming known ones [45]. Also, a deep convolutional neural network was used to monitor the electromagnetic emissions of a QKD emitter (Deep-learning-based radio-frequency side-channel attack [46]). Despite neural networks demonstrating a better ability to perform quantum state reconstruction using partial information and fewer measurements than classical state tomography [47, 48, 49, 50], the literature does not show extensive research on ML/DL-based attacks.
@@
In this paper, we develop a new type of individual attack based on continuous measurement [51] of single polarized photons and apply it for concreteness in the context of the BB84 protocol, by contrast to other works that usually apply this kind of measurement on CVQKD. The general motivation is to evaluate the performance of attacks that produce only a small perturbation to the qubits, by contrast with the effects of projective measurement usually involved in the other types of attacks. In particular, we investigate how the effects of the spy measurement can be optimally hidden by the intrinsic noise of the quantum communication channel by minimizing the increase in QBER due to the measurement. At the same time, we investigate how neural networks can effectively use the information extracted by these types of measurement, to infer a more significant part of the key than conventional means. To do so, we feed the outcome of the continuous measurement, also called homodyne photo currents, to a Long Short-Term Memory (LSTM) recurrent neural network [52] to retrieve the initial states of the photons sent by Alice, which compose the sifted key generated. The main point of this paper is not to question the security of the BB84 protocol, but rather to show an application of deep learning in QKD, and to raise awareness on its use in the context of attacks.
@@
This paper is organized as follows: In Sec. II we first summarize the BB84 protocol, present our model of the qubit dynamics in the quantum communication channel when subject to intrinsic dissipation and continuous measurement, investigate how a spy could use the outcome of this measurement to obtain the initial state of the qubit and present the neural network we implemented to do so. In Sec. III we compare the results obtained via a basic projective measurement (Intercept-and-Resend attack) and our measurement scheme. We then extend the theory to include quantum feedback, to try to decrease the impact of the measurement while keeping the same accuracy of the neural network. In Sec. IV., we discuss our attack scheme in terms of information gain and how it fits into the thresholds established by the information-disturbance principle, and calculate the typical key rate Alice and Bob should achieve to secure the protocol. In particular, we compare our attack performances against the ones of an optimal individual attack strategy for the BB84 protocol: the covariant-phase quantum cloner [53, 54]. Finally, in Sec. V, we conclude and discuss potential perspectives of our work.
@@ II Model and methods
In this section, we first remind how the standard BB84 protocol works briefly, before presenting how we model the dynamics of the qubits used in the protocol when they are subject to dissipation and continuous measurement in the quantum communication channel. Then, we describe the neural network that we envision a spy could use to retrieve the states of the qubits based on the continuous measurement they performed in the channel. Finally, we present how we quantify the impact of the measurement on the protocol.
@@ 5. Step 5:
    Alice selects a subset of her bits to check on the interference caused by the spy – the so-called Eve –, and tells Bob which bits she chose. They both announce and compare the values of the check bits via the public channel and calculate the QBER given by Eq. (1). If it is higher than a threshold (typically $11\%$ [30]), they abort the protocol.
Alice and Bob now each possess a sifted key, which may slightly be different because of the dissipation and spy-induced QBER. To increase the security of the protocol, two additional steps are performed. The first is information reconciliation (also called error correction) to correct bits that have been modified by dissipation and spying [56, 57]. The second is privacy amplification, which consists in passing the generated key in a hash function to exponentially decrease Eve information. [56, 58, 59]. By doing so, Alice and Bob obtain a shorter but more secure final key. It is important to note that hash functions decrease the key length, such that the spy looses information in the data compression process. This last step requires, however, an upper bound estimate of Eve’s information on the corrected key. Thus the bigger the part of the sifted key Eve has, the more bits Alice and Bob must sacrifice in the process.
@@ II.2 Dissipative qubit dynamics conditioned on measurement
Figure 1: Sketch representation of our attack scheme applied on the BB84 protocol. Alice sends linearly polarized photons to Bob via an optical fiber, while traveling they are subject to intrinsic dissipation in the fiber with a rate $\gamma_{D}$ and to weak measurement with a rate $\gamma_{E}$ performed by Eve on a certain portion of the fiber, the output of which being treated by a neural network to determine the initial state. In the sketched example, Alice sends six photons to Bob, among which three are polarized in the Pauli-X eigenbasis and three in the Pauli-Z eigenbasis. Bob measures randomly in one of these two bases each of the photons received, and the basis chosen is the right one for four of these. However, even if the measurement basis is right, they are two errors. The sketch highlights the two possible mechanisms for errors: intrinsic dissipation (blue dashed square) and measurement by a third party (purple dashed square). In this example, if Alice and Bob compare all the measurements when they chose the same basis, the QBER [Eq. (1)] would be $50\%$.
@@
$+\gamma_{E}\mathcal{D}[e]\rho_{J}dt+\sqrt{\gamma_{E}\eta}\mathcal{H}[e]\rho_{J}dW,$
where $H$ is the channel Hamiltonian defined as $H=\omega\sigma_{z}$ for the initial states |$\ket{0}$ or |$\ket{1}$ and $H=\omega\sigma_{x}$ for $\ket{+}$ or $\ket{-}$ [26] with $\sigma_{x}=|0\rangle\langle 1|+|1\rangle\langle 0|$ and $\sigma_{z}=|0\rangle\langle 0|-|1\rangle\langle 1|$ the standard Pauli operators, where $\rho_{J}$ is the density operator of the qubit conditioned on the measurement with efficiency $\eta\in[0,1]$ of the homodyne current [60]
@@
to model the dissipation as a bit-flip error, but any other choice could be made without any additional complexity (see Sec. IV.2), depending e.g. on the specific open system model considered for the optical fiber.
In terms of practical implementations of the continuous measurement, while we do not intend in this paper to provide a specific detailed scheme, we foresee that this could be realized indirectly, for examples, via the monitoring of an auxiliary field that couples to the photons on a certain portion of the optical fiber, or by extracting a small amplitude of the signal via a low-ratio beam splitter or a directional coupler. In Appendix A, we show how to derive Eq. (3) from the homodyne detection of such an ancillary field and its adiabatic elimination.
@@ II.4 Neural network quantum tomography based on the measurement
The input layer of our network is a LSTM one with $40$ units in its hidden state, to take as input the time series data that are the photo currents. We use the last hidden state (dimension $40$) as the input of a $40$-neurons dense layer, followed by another $20$-neurons dense layer, with activation functions set to ReLU and sigmoid respectively. The output layer is a $4$ neurons softmax layer. The loss function we use for training is the sparse categorical cross entropy since we deal with a $4$-class classification problem, and the optimizer is Adam with default parameters. The training consists in one epoch over $9\times 10^{4}$ photo currents while the testing is done on $10^{4}$ photo currents. Before entering the network the currents are standardized and flipped along their time axis.
@@
II.5 Impact of measurement VS Accuracy
As explained earlier, the goal of the spy is to minimize their impact on the photon states while maximizing the part of the sifted key obtained. To quantify the impact of the measurement, we use the QBER introduced earlier [Eq. (1)], as it is directly measurable by Alice and Bob and allows them to assess the security of the protocol. To quantify the success of the spy in retrieving the initial state of the photons, we use the spy accuracy that we will denote by $A$. In the context of our deep learning approach, this is the neural network test accuracy [72] which is defined as the percentage of good predictions among all the predictions of the network on the test set. However, when dealing with a projective measurement, this accuracy is defined as the probability that the spy measures the right state. Note that the amount of information extracted from the qubits (e.g., the information gain) is defined by the measurement scheme. On the other hand, accuracy $A$ depends on both the information gain and the ability of the neural network to efficiently harness it, as detailed in Sec. IV.1. Thus $A$ does not correspond properly to a measure of the extracted information in the sense of Shannon [73]. However, we use it for convenience to quantify the success of the eavesdropping scheme, since it is a performance measure that represent the average percentage of the sifted key obtained by the spy.
@@ III Results
In this section, we study the impact of our attack and its performance in different cases in terms of QBER and accuracy $A$. We first compute the QBER in the case of no attack. Then, we study a simple standard projective measurement attack, before investigating our continuous measurement scheme. We also study the performance of quantum feedback the spy could use to cover their tracks. Note that in this section, except stated otherwise, all time durations are measured in units of $1/\gamma_{D}$.
@@ III.1 No attack
In the case where no spying is done on the quantum channel, which means there is no measurement and only the intrinsic dissipation, the QBER can easily be obtained from Eq. (3) with $e=0$, which corresponds to a Lindblad master equation (see Appendix B), and reads
@@ III.2 Attack via projective measurement
In this case, the accuracy $A$, can be calculated exactly by solving the Lindblad master equation with a single jump operator $L=\sigma_{x}$ (see Appendix C) and reads
@@
which depends on the time $t^{*}$ at which the projective measurement is performed. Therefore, Eve must measure the photons as close to Alice as possible in order to maximize Eq. (9) and get as much as possible of the sifted key, which is here bounded by $75\%$, meaning that the spy has at best $75\%$ chance to guess the initial state sent by Alice.
The QBER can also be obtained easily (see Appendix D), and reads
@@
Interestingly, we see that the time $t^{*}$ at which Eve performs her measurement does not impact the probability that Bob measures the state he is supposed to. Also, comparing Eqs. (8) and (10), we clearly see that Alice and Bob will easily distinguish the presence of the spy from intrinsic dissipation. This result is displayed as the dotted black line in Fig 3.
Figure 3: Top: QBER as a function of $\gamma_{D}t$ for different attack schemes and angles $\theta$. The black dashed line represents the QBER when no measurement is performed (i.e., produced only by intrinsic dissipation) [Eq. (8)]. The dotted black line is the QBER with dissipation and projective measurement [Eq. (10)] and the red line is the QBER caused by a continuous measurement with $e=\sigma_{z}$ ($\theta=\pi/2$). The orange line is the QBER when using the angle $\theta_{A}^{*}$ maximizing the accuracy (which we update every $\gamma_{D}t=0.3$), the green line for the angle $\theta_{\mathrm{qber}}^{*}$ minimizing the QBER, and the blue line for the angle $\theta_{\lambda}^{*}$ minimizing $\lambda$. Bottom: optimized angles with respect to the QBER itself (green line), the NN test accuracy (orange line) and $\lambda$ (blue line).